# LLMs’ Reshaping of People, Processes, Products, and Society in Software Development: A Comprehensive Exploration with Early Adopters

### v1

@@ 3.4.2. Education
  • Needing to Adapt: Six participants acknowledged the flip side of the impracticality of banning LLM usage, suggesting ways that computing education programs should adapt to the new reality with LLMs. Four participants expressed that the ease of plagiarism with LLMs encourages students to use them for ready-made solutions, introducing a need for change. Two felt cheating was going to happen, with P16 suggesting that cheating is inherent to certain individuals: “People who cheat, they’re going to cheat anyway, and that’s gonna get caught up with them. That’s a character trait.” P8 mentioned, “I think [students] don’t write any code in their homework based on the [access to] ChatGPT because they already have whatever they want.
    Two participants suggested that instructors mitigate the impact of cheating concerns by redesigning assignments to ensure they require a genuine understanding of concepts through critical thinking and problem-solving skills. P13 suggested comparing student work to LLM answers: “The professors would have to run it through GPT to see how it’s performing, and see what kind of variations of answers it’s generating.” P2 additionally shared, “You can design the assignments in such a way that even if [students] use [ChatGPT], they need to understand the concepts.
    Two more participants mentioned the importance of preparing students for coding without LLM assistance. Although P5 was generally a proponent of integrating ChatGPT into the curriculum, he stressed the importance of independence: “You need to be able to create code on your own, but also collaborate with others as well.
4. Discussions
As outlined in Section 3, we have highlighted the advantages and challenges of current LLMs across four key dimensions. Based on our findings:
4.1. Implications for Future Software Developers
4.1.1. Educating future software developers on prompt engineering
Our participants emphasized the importance of prompt engineering as a key technique for effectively utilizing LLMs. This underscores the need for workshops and training sessions to focus on prompt engineering, particularly on strategies such as maintaining specificity and brevity, formulating problems clearly, and considering linguistic nuances for optimal results.
To support this learning, developers can access formal resources, such as OpenAI’s comprehensive guide on prompt engineering (OpenAI, [n. d.]) or DeepLearning.AI’s free [^10] At the time of writing. course on prompt engineering practices for developers (Fulford and Ng, 2024). These resources provide valuable insights and practical techniques to enhance their expertise in working with LLMs.
4.1.2. Educating on problem decomposition
The ability to break down complex problems into manageable components is crucial for maximizing the effectiveness of LLMs. Our findings indicate that LLMs perform optimally when presented with tasks that are clear, concise, and limited in scope. By developing strong problem decomposition skills, developers can better align their challenges with the capabilities of LLMs, significantly improving the efficiency and accuracy of their solutions. Mastering this skill empowers developers to leverage LLMs as powerful tools in addressing a wide range of software development tasks.
4.1.3. Setting realistic expectations for LLM use
Developers must recognize the inherent limitations of LLMs, including issues such as unreliable responses (Xu et al., 2022), hallucinations, knowledge gaps, and difficulties in maintaining consistent contextual understanding (Hou et al., 2024). Although LLM accuracy and effectiveness continue to improve, their performance remains variable across different domains. Understanding which domains to trust and the extent of reliability is essential for using these tools effectively.
One practical strategy is to explicitly identify the software engineering tasks for which LLMs excel or underperform, as outlined in this paper (see Section 3). Additionally, equipping developers with a deeper understanding of how LLMs function can bridge gaps in expectations and capabilities. Structured courses or workshops tailored to software developers’ needs could offer a robust foundation, similar to existing workshops designed for educators(Reichert et al., 2024).
These educational initiatives not only enhance developers’ ability to use LLMs effectively but also provide the added benefit of deepening their technical knowledge, enabling them to adapt to and capitalize on future advancements in generative AI.
4.1.4. Adopting LLMs as Surrogate Team Members
Freelance developers, or those who primarily work independently, encounter a distinct set of challenges compared to their counterparts in collaborative environments. Insights from two study participants (P9 and P10) indicate that LLMs provide specific advantages uniquely suited to solo developers, which may not be as pronounced for other developer types. Based on our findings, freelance or solo developers can utilize LLMs as surrogate team members, serving as virtual colleagues, subcontractors, or assistants. LLMs can support these developers by refining and clarifying requirements, conducting research, discovering information during stakeholder interactions, and simulating code reviews to uphold professional coding standards. Moreover, LLMs can reduce—though not entirely eliminate—the need for subcontractors, thereby enhancing the efficiency and self-sufficiency of independent developers.
4.1.5. Encouraging effective use of LLMs for programming
Despite their limitations, LLMs offer considerable advantages to developers who understand how to harness their capabilities effectively. As indicated by our findings, developers can use LLMs to enhance various aspects of their programming practices, including: (1) Refactoring Code: Improving the structure and organization of existing code for maintainability and performance. (2) Learning and Applying Design Patterns: Gaining insights into widely recognized solutions to common programming challenges. (3) Enhancing Code Readability: Producing more understandable and clean code, which facilitates collaboration and long-term project sustainability. (4) Automating Boilerplate and Repetitive Tasks: Quickly generating routine code components to save time and focus on more complex challenges. These benefits not only streamline workflows but also provide opportunities for developers to adopt and reinforce best practices. Our findings align with similar research (e.g., (Mendes et al., 2024; Li et al., 2024; Pinto et al., 2024; Coutinho et al., 2024) highlighting how LLMs can optimize developer productivity and programming quality. By using LLMs strategically, developers can maximize their efficiency and improve both individual and team outcomes.
4.1.6. Emphasizing secure practices
Developers must remain vigilant about the proprietary and security aspects of the code they share with LLMs, as emphasized by Wang et al. (Wang et al., 2023). Inputted data such as code, prompts, and other personal information may be collected and utilized by companies like OpenAI and Gemini to improve their services [^11] At the time of writing, both OpenAI (n.a., 2024d) and Gemini (n.a., 2024c) state that they collect and use personal data, including prompts and log data, as the default setting to improve their model and services.. This makes it crucial for developers to ensure compliance with contractual agreements and address licensing issues to minimize risks when using these tools.
Another critical concern is that LLMs often do not provide the source of the code they generate (Dou et al., 2024) [^12] Pre-published via arXiv., leaving the origins and security of the produced code uncertain. This ambiguity poses potential risks, particularly in production environments or sensitive projects. Developers engaged in research or ideation may perceive fewer security challenges, but adhering to robust security protocols is essential for all.

### v2

@@ 4.4.2 Education
  • Needing to Adapt: Six participants acknowledged the flip side of the impracticality of banning LLM usage, suggesting ways that computing education programs should adapt to the new reality with LLMs. Four participants expressed that the ease of plagiarism with LLMs encourages students to use them for ready-made solutions, introducing a need for change. Two participants suggested that instructors mitigate the impact of cheating concerns by redesigning assignments to ensure they require a genuine understanding of concepts through critical thinking and problem-solving skills. P2 shared, “You can design the assignments in such a way that even if [students] use [ChatGPT], they need to understand the concepts.” Two more participants mentioned the importance of preparing students for coding without LLM assistance.
5 Discussion
5.1 The Sociotechnical Transformation of Software Development with LLMs
Our investigation of early adopter experiences reveals that LLM integration in software development represents a fundamental sociotechnical transformation, rather than simply adopting a tool. Across our four research dimensions—people, process, product, and society—we observe a consistent pattern: LLMs provide substantial benefits when developers possess specific competencies, work within supportive organizational contexts, and maintain critical evaluation practices. The technology itself is neither inherently beneficial nor harmful; rather, its impact depends on the sociotechnical system within which it operates. This finding aligns with established sociotechnical systems theory in software engineering, which emphasizes that technology adoption succeeds or fails based on the interplay between technical capabilities, human practices, organizational structures, and broader social contexts [71].
These dimensions are deeply interconnected. Developers’ growing confidence and productivity (RQ1) enabled more sophisticated process adaptations (RQ2), but only when coupled with rigorous verification practices that maintained product quality (RQ3). Notably, we observe a productivity-quality paradox: participants reported significant time savings yet frequently discarded generated code (eight participants discarded more than 50%, two discarded approximately half). This apparent contradiction resolves when we recognize that LLMs fundamentally shift developer effort from generation to evaluation—a transformation from syntactic production to critical assessment that requires a deep understanding to identify subtle errors, architectural mismatches, and security vulnerabilities. This shift has profound implications for developer expertise, organizational workflow, and educational preparation (RQ4), as the profession grapples with redefining what it means to “write code” versus “orchestrate code generation.”
Our findings illuminate three critical themes that transcend individual research questions. First, effective LLM use requires new competencies that differ fundamentally from traditional programming skills—prompt engineering, critical evaluation, and security-conscious integration (detailed in Section 4.2.3). Second, LLM utility is highly phase-dependent across the software development lifecycle, with near-universal adoption during implementation but minimal impact on requirements gathering (examined in Section 4.2.1). Third, human judgment, collaboration, and domain expertise remain persistently important despite automation of certain tasks. Six participants explicitly noted LLMs’ inability to replace human interactions, twelve emphasized that requirements gathering remains fundamentally human-centered, and eight stressed the critical importance of problem decomposition skills that LLMs cannot provide.
These themes underscore that successful LLM integration requires not just technological sophistication but also organizational support, educational reform, and ongoing professional development. We explore each theme in depth below, positioning our findings within the broader empirical literature and deriving practical implications for multiple stakeholder groups.
5.1.1 The Productivity-Quality Paradox
As introduced in Section 4.1, the productivity-quality paradox, i.e., developers reporting substantial time savings while discarding much generated code, challenges simplistic narratives about LLM productivity and requires careful interpretation. This pattern reveals that productivity manifests through multiple mechanisms rather than direct code reuse.
LLMs shift developer effort from code generation to evaluation and integration. Productivity emerges through: (1) rapid prototyping and iteration, where developers quickly explore multiple approaches even if individual attempts are discarded, as P4 explained: “I’ve kind of isolated its use cases to helping me improve code at the function level, like small snippets of code, and helping me go from zero to something”; (2) learning through examples, where generated code serves as educational material—eleven participants highlighted LLMs’ ability to explain code and demonstrate best practices; (3) overcoming cognitive blocks, where LLMs help developers get “unstuck” by suggesting approaches or providing boilerplate; and (4) reducing search friction, where LLMs aggregate information from multiple sources, as P14 noted: “I don’t need to click on each of the links from Google to find the solution […] [ChatGPT] gives me a summarized version from possibly three, four different sources.”
This reframing reveals LLMs’ role as thinking partners or ideation tools rather than code factories. The value lies in cognitive support, i.e., helping developers think through problems, explore solution spaces, and learn approaches, rather than in generating production-ready code. Fifteen participants modified generated code before integration, and twelve copy-pasted as a starting point for adaptation rather than final implementation. As P12 articulated: “You can take [Bard’s response] as an idea, like a first draft, and then iterate on it; that’s the experience I have working with it from the coding side.”
However, this productivity shift introduces new risks. The ease of generating plausible-looking code may reduce cognitive engagement necessary for deep learning and quality assurance. Five participants reported becoming more skeptical in their code review practices specifically because of LLM-generated code, with P1 noting: “[I]n a sense, yes, it has changed testing. I’m a bit more skeptical.” This heightened skepticism is appropriate. The productivity-quality paradox thus reveals both opportunity, i.e., LLMs can accelerate development when used strategically and responsibility, i.e., developers must cultivate stronger evaluation skills to ensure productivity gains do not compromise code quality, security, or learning.
5.2 Positioning Our Findings: Early Adoption Patterns and Subsequent Validation
Our findings as early-adopter data collected in mid-2023 provide a valuable lens into emergent patterns that subsequent large-scale research has also reported. The convergence between our qualitative insights and subsequent quantitative studies demonstrates both the reliability of our findings and the value of early adopter perspectives in understanding technological adoption trajectories.
Several key findings have been corroborated by large-scale industry research conducted between 2024 and 2025. Our participants’ reports of substantial productivity benefits align with GitHub and Accenture’s 2024 study finding that 96% of developers report faster task completion with AI tools, with an average productivity increase of 55% [12], further corroborated by a systematic literature review carried out by Hou et al [25]. The security concerns we documented align with Stack Overflow’s 2024 developer survey, which identifies security as the top concern about AI tools, with 45% of developers worried about data privacy implications [11], further corroborated by Islam. et al in their 2024 work on LLM induced vulnerabilities [43]. Our finding that LLMs facilitate learning is echoed in recent educational research, which demonstrates LLMs’ effectiveness as personalized learning tools while raising concerns about potential learning impediments [14]. The prompt engineering strategies we documented have since been formalized in training materials and research on effective prompting patterns [24].
This convergence validates our early-adopter methodology: patterns observed among pioneers who adopted LLMs between November 2022 and April 2023 proved predictive of broader adoption trends visible in studies with thousands of participants conducted 12 to 18 months later. Early adopters serve as “scouts” for technological change, developing practices and encountering challenges before they become widespread. Innovation diffusion theory identifies them as typically more educated, higher in social status, and more willing to tolerate uncertainty [72]. Our participants exemplify these characteristics, developing sophisticated practices through intensive experimentation that later adopters can now learn through structured training.
However, convergence alone understates our contribution. Our qualitative approach and comprehensive SDLC focus reveal insights less visible in survey-based or tool-specific studies, making three distinctive contributions to the literature. First, our phase-dependent adoption analysis, showing minimal requirements-gathering impact (12/16) despite near-universal implementation adoption (16/16), provides granular insight into where and why LLMs prove useful across the development lifecycle. Most existing studies focus narrowly on coding tasks [73] or specific tools [41], while our comprehensive SDLC mapping reveals that LLM utility is highly context-dependent, challenging oversimplified narratives about LLMs “transforming software development.”
Second, our taxonomy of six prompt engineering strategies and five verification methods, derived from observed practice rather than prescribed techniques, provides empirical grounding for understanding how developers actually use LLMs. Our early documentation captures the emergent, experimental nature of these practices before they became formalized, revealing not just what works but how developers discovered effective approaches through trial and error.
Third, our finding that high code discard rates (8/16 discard ¿50%) coexist with productivity gains challenges simplistic “productivity improvement” metrics prevalent in industry reports. This productivity-quality paradox reveals that LLMs’ value lies in cognitive support (ideation, exploration, learning) rather than direct code reuse. This nuance, which requires a qualitative investigation to uncover, fundamentally reframes how we should measure and understand LLM “productivity.” Survey-based studies asking “Do you find LLMs productive?” capture perceived value but miss the complex reality of substantial code rejection alongside productivity benefits.
Our study’s distinctive value emerges from the intersection of three methodological choices: (1) timing, capturing early adoption before practices became standardized, (2) method, qualitative depth revealing nuances invisible in surveys, and (3) scope, comprehensive SDLC coverage rather than narrow task focus. By documenting early-adopter experiences during the formation stage of LLM integration practices, our study captures not just outcomes but processes, the reasoning, experimentation, and sense-making through which effective practices emerged. This developmental perspective is valuable for supporting later adopters, who benefit from knowing not only what to do but also why certain approaches work and how to adapt them to specific contexts.
5.2.1 For Individual Developers: New Competencies and Practices
Effective LLM integration requires developers to cultivate competencies that differ from traditional programming skills. Our findings reveal three critical capability clusters: (1) prompt engineering and problem formulation, (2) critical evaluation and verification, and (3) security-conscious integration. These competencies complement rather than replace traditional skills; developers still need strong fundamentals in data structures, algorithms, and software architecture to evaluate outputs, identify errors, and make sound design decisions. As P5 emphasized, LLMs require “a solid foundation to pose the right queries.”
Our findings provide substantial clarification to a distinct, role-based divergence in both LLM application and perceived risk. We found that junior developers, in particular, derived significant educational value; all junior participants (4/4) felt that access to an LLM helped foster their learning; two framed the LLM as a “pair-programmer” (P5, P9) that accelerated their onboarding and understanding of unfamiliar code. Conversely, senior developers primarily leveraged LLMs to streamline software development workflows by offloading mundane tasks, thereby preserving their cognitive resources for complex architectural challenges. However, our study also complicates this narrative by identifying a tension in perceived risk. While junior developers focused on the immediate learning benefits, it was predominantly senior developers (5/8 participants in this theme) who articulated fears of stagnation of skill development, expressing explicit concern that over-reliance could atrophy the foundational problem-solving skills of their less-experienced colleagues. Our work thus refines the ”role-dependent” hypothesis by specifying how use cases diverge and revealing that concerns over skill erosion are themselves stratified by seniority.
Prompt Engineering and Problem Formulation
Prompt engineering emerged as a sophisticated practice encompassing six distinct strategies. Fifteen participants varied the specificity of prompts based on desired outputs, recognizing that broad prompts encouraged creative solutions, while specific prompts yielded targeted answers. As P15 explained: “[I]t started to give me actually better results, because it knew things that I didn’t know […] When I have a more broad [prompt], then it’s able to kind of formulate its own deduction.” Conversely, P2 advocated specificity for software tasks: “I think it should be very specific […] for software development, I think you need to be very specific.” The strategic lesson is that effective prompters adapt their specificity to context.
Beyond specificity, participants employed iterative refinement (fifteen used follow-up queries), security-conscious generalization (thirteen avoided sharing confidential information), contextual examples (nine practiced few-shot prompting), conversation management (nine started new threads when context degraded), and deliberate practice (fourteen acknowledged prompt engineering as essential). As P16 noted: “The follow-up questions actually might be more important than [the initial question].”
These strategies require metacognitive awareness; understanding not just the problem but also how to formulate it for an LLM’s particular strengths and limitations. This represents a shift from directly solving problems to orchestrating problem-solving, a form of meta-programming that complements traditional coding skills.
Closely related is problem decomposition. Eight participants noted that LLMs work best with small, well-defined tasks, necessitating more deliberate decomposition. Interestingly, some viewed this as a “forcing function” encouraging better practices. P4 shared: “[ChatGPT] might force me to break [problems] down […] If I can’t explain it to ChatGPT, then it might be an indicator that I don’t yet understand it.” This suggests that LLM constraints may inadvertently reinforce fundamental SE principles.
Developers can build prompt engineering competencies through formal resources such as OpenAI’s comprehensive guide [74] or DeepLearning.AI’s course [75], though experiential learning, i.e., deliberately practicing on non-critical tasks and reflecting on what works, is equally valuable for developing the metacognitive awareness that characterizes expert LLM use.
Critical Evaluation and Verification
Our participants employed five verification methods to ensure code quality: reading generated code (ten participants), checking outputs against expectations (ten), manual testing using developer tools or REPL environments (seven), external testing frameworks (four), and consulting external sources, such as documentation (two). These practices reveal a multi-layered verification approach where developers combine automated and manual techniques to build confidence in generated code.
Five participants reported becoming more skeptical and vigilant in their code review practices, particularly due to LLM-generated code. As P6 advised: “Be wary of what you get […] Don’t give up on checking the responses.” This heightened critical stance is appropriate: LLM-generated code can appear syntactically correct while harboring subtle logical errors, using outdated APIs, or introducing security vulnerabilities. Two participants explicitly stated that LLM code should not be copied without review, with P11 noting: “I think it’s dangerous, right? Like, if out of our work on a team, somebody would just copy-paste ChatGPT [code], you know, I’ll probably be annoyed by it.”
The verification burden explains our productivity-quality paradox: developers save time during generation but must invest substantial effort in evaluation. Fifteen participants modified code before integration, making substantive adaptations to fit their specific contexts, correcting errors, and aligning with project standards. This pattern reveals that LLM’s “productivity” manifests through rapid iteration and exploration rather than direct code reuse, requiring developers to possess the expertise necessary to critically assess and adapt generated outputs.
Security-Conscious Integration
Security-conscious integration emerged as a critical practice, with participants demonstrating bidirectional awareness: protecting proprietary data when prompting and scrutinizing generated code for vulnerabilities. Thirteen participants generalized prompts or used case-specific information rather than project-specific details. As P2 explained: “Because I cannot put all the information into ChatGPT, because it’s very client-specific information, it’s confidential information […] I just give it use cases that are similar.”
This data protection practice reflects appropriate caution: code and prompts inputted into LLM tools may be collected and used by companies to improve their services. Developers must ensure contract compliance and address licensing issues to mitigate risks. Some participants employed additional protections; P14 used a VPN, while P4 and P15 relied on careful code review as a security layer, but these individual measures cannot substitute for clear organizational policies.
On the output side, eight participants expressed concerns about using LLM-generated code without thorough review. Since LLMs often do not provide the source of generated code [76], the origins and security remain unclear, particularly concerning for production environments. P1 expressed concerns about malicious actors potentially injecting poisoning code that could propagate through training, i.e., a supply chain security concern extending beyond individual developer practices.
Thirteen participants considered LLMs’ code sufficiently secure for their purposes, but this confidence was context-dependent: participants primarily used LLMs for learning, prototyping, or non-production development. As P2 noted: “I’ve never had to use it for any security aspects.” This underscores that security practices must be calibrated to use cases, with more stringent standards for production code.